# From Code to Play: Benchmarking Program Search for Games Using Large Language Models

### v1

@@ V-B Vehicle Driving
Llama 3.1 405B 145.84 130.82 109.03 105.64 80.97 87.46 57.69 64.41 82.57 57.11 92.15
Llama 3.1 8B 186.14 186.14 186.14 177.24 116.08 186.14 186.14 131.11 106.02 179.77 164.09
Llama 3.1 70B 111.84 132.64 147.98 111.67 80.97 64.53 79.61 106.37 74.42 104.63 101.47
Claude 3.5 Haiku 174.29 142.54 95.95 86.20 102.06 76.60 103.84 76.42 93.45 67.51 101.89
Claude 3.5 Sonnet 153.04 122.59 102.66 88.09 83.57 85.54 78.21 85.75 66.23 78.59 94.43
GPT 4o mini 154.57 136.83 118.76 128.51 102.21 110.48 111.68 111.30 104.90 97.49 117.67
GPT 4o 137.02 142.58 111.51 111.69 104.16 105.40 96.30 85.88 57.98 94.43 104.70
o1-mini 130.86 167.24 114.41 111.67 105.38 96.21 68.51 85.79 93.41 94.43 106.79
Mistral Small 186.14 186.14 186.14 180.56 120.38 117.30 108.41 110.75 104.90 186.14 148.69
Mistral Large 186.85 121.88 96.40 83.88 101.07 68.66 97.34 58.75 92.38 94.76 100.20
Gemini Flash 156.79 145.74 149.34 178.73 111.47 110.73 115.08 117.94 186.14 186.14 145.81
Gemini Pro 180.41 122.72 146.10 75.36 106.88 105.51 112.58 154.27 102.75 114.61 122.12
@@
Table IV shows the results of the driving experiments for different rotation speeds $\omega$ to adjust the difficulty level of steering the asteroid ship. The numbers are the minimum distance achieved by the best program for five evaluation episodes. $D_{avg}$ is the average of all distances for each LLM. The experiments were conducted with 10 iterations of search and three attempts to generate or improve the program. We omitted the number of successful iterations because no LLM managed to stop the asteroid ship at the target position in all five evaluation episodes. A program was considered successful only if it could consistently stop the vehicle within a specified tolerance $t$ across all evaluation episodes. In our experiments, the synthesized programs succeeded in stopping the vehicle in only one or two episodes within a tolerance of $t=10$, and thus no program qualified as successful. The overall best model is Llama 3.1 405B followed by Claude Sonnet. Regarding $D_{avg}$ larger models generally outperformed their smaller counterparts, with o1-mini being slightly worse than GPT-4o. For the different rotation speeds, Llama 3.1 405B and 70B generated the best program in three out of 10 tasks, for the other tasks the best programs were synthesized by the large LLMs, except for $\omega=30$, where Claude Haiku found the best program. As no LLM successfully solved the problem with a simple prompting strategy in this experiment, we consider it a compelling challenge for future research.
@@
TABLE V: Highest reward per language model and level (with number of successful runs per level).
@@
Llama 3.1 8B 95 (7) 0.0 (0) 0.0 (0) 95 (5) 0.0 (0) 0.0 (0) 58 (2) 95 (5) 0.0 (0) 0.0 (0) 4
Llama 3.1 70B 95 (7) 0.0 (0) 0.0 (0) 0.0 (0) 0.0 (0) 0.0 (0) 94 (3) 95 (3) 0.0 (0) 0.0 (0) 3
Llama 3.1 405B 95 (9) 0.0 (0) 0.0 (0) 95 (1) 0.0 (0) 0.0 (0) 94 (7) 95 (4) 0.0 (0) 90 (1) 5
Claude 3.5 Haiku 95 (10) 0.0 (0) 0.0 (0) 95 (5) 0.0 (0) 25 (1) 94 (8) 95 (10) 0.0 (0) 0.0 (0) 5
Claude 3.5 Sonnet 95 (10) 0.0 (0) 0.0 (0) 95 (9) 0.0 (0) 91 (2) 94 (6) 95 (9) 90 (2) 92 (1) 7
@@
GPT 4o 95 (10) 0.0 (0) 0.0 (0) 95 (4) 0.0 (0) 91 (1) 94 (5) 95 (8) 0.0 (0) 0.0 (0) 5
o1-mini 95 (10) 0.0 (0) 0.0 (0) 95 (7) 0.0 (0) 83 (1) 94 (7) 95 (9) 0.0 (0) 92 (1) 6
Mistral Small 95 (5) 0.0 (0) 0.0 (0) 95 (3) 0.0 (0) 0.0 (0) 33 (1) 95 (5) 0.0 (0) 0.0 (0) 4
Mistral Large 95 (10) 89 (1) 0.0 (0) 95 (1) 0.0 (0) 0.0 (0) 94 (8) 95 (9) 0.0 (0) 0.0 (0) 5
Gemini Flash 95 (10) 0.0 (0) 0.0 (0) 0.0 (0) 0.0 (0) 83 (1) 94 (5) 95 (8) 0.0 (0) 90 (1) 5
Gemini Pro 95 (6) 0.0 (0) 0.0 (0) 0.0 (0) 0.0 (0) 83 (4) 76 (5) 95 (6) 0.0 (0) 0.0 (0) 4
V-C Baba is You
@@
For our experiments, we used a Python version [^11] https://github.com/ADockhorn/Keke-AI-PY of the Keke is You AI framework [18]. For this domain, we prompted the LLMs to provide a policy, giving a short description of the game and the initial state of the level (the complete prompt is given in the repository). Similar to the Minatar experiments, the state is converted into a text description. The function to be written should use the current state as input and return a movement direction or the command for waiting a turn. Each episode ends after 100 actions or once the win condition is fulfilled. A reward is awarded based on the maximum number of actions (100) minus the number of steps taken. Thus, the return can be maximized by finishing the level as fast as possible. Each level can be solved in less than 20 actions.
In our tests, we queried the agent to solve 10 simple demo levels (see Figure 4). Each of the levels focuses on one or multiple key mechanics of the framework such as rule interpretation (levels 1-10), rule creation (levels 2, 3, 5) or destruction (levels 6, 8, 9, 10), and object manipulation (level 7). Table V shows the results of our comparison of the LLM models’ capabilities. Each entry shows the reward of the best program after 10 iterations, in which each iteration used 3 queries to generate or improve the policy. The number of successful iterations is shown in brackets.
Most agents were able to solve at least 5 out of 10 levels, with Claude 3.5 Sonnet being the only model able to solve 7. For Claude, GPT and Mistral, models of the same vendor with higher number of parameters were able to solve more levels. For the Llama 3 models, the 8B model solved one more level than the 70B model but was outperformed by the 405B model. Similarly, Gemini Flash performed slightly better than Gemini Pro. Tested models were mostly successful in interpreting existing rules. As can be seen, some levels are rarely solved by any model. Creating or destroying rules and thus modifying the logic of our game world has proven difficult for all models. Nearly all models have failed in solving levels 2 and 3, which require rule creation, and levels 9 and 10 which require a rule’s destruction to finish the puzzle. Slight differences in the observed success rate could be due to the low number of repetitions per level resulting in sampling errors: levels 2, 9, and 10, which are rarely solved at all, could be affected by this. Chain of thought prompting [49] may potentially help in overcoming these more complex planning tasks.
V-D Procedural Content Generation
Figure 5: Three generated mazes by algorithms aiming to optimise the longest shortest path objective. Left: (score 18) example from a simple LLM generated algorithm setting wall cells with a fixed probability. Middle: (score 38) example a more sophisticated LLM algorithm involving recursion and a shortest path algorithm. Right: (score 54) example from an evolutionary algorithm directly optimising for the objective function.
@@ V-D Procedural Content Generation
The prompt advised the LLMs to use the longest shortest path objective to guide the maze generation process. This objective encourages intricate and interesting mazes. Most of the generated code ignored the hint and instead coded overly simple algorithms, placing corridors and walls in each cell with a given probability while usually ensuring that the start and end points were not on wall cells. An example generated maze is shown in the left of figure 5. Occasionally, a better algorithm was produced that mixed randomness, recursion and graph search in ways we’ve not fully analysed. These algorithms sometimes produced mazes with no path between start and end, resulting in a score of -1. When they worked, they often produced reasonable mazes such as the one shown in the middle of Figure 5. The LLMs failed to find an algorithm as effective as an evolutionary algorithm applied to directly solve the objective. A sample maze from such an algorithm is shown on the right of the figure.
@@
Note that here we are evaluating the effectiveness of the algorithms in meeting the specified objective, which is to produce mazes with the longest shortest path between start and end. Depending on the application, this could be a poor objective to maximise, with the best mazes have a mid-ranking score, such as the central maze in Figure 5.
@@
TABLE VI: Maze generation LLM results. In each case 10 iterations of search were run, with up to three attempts to improve or generate the policy. $D_{max}$ is the maximum distance of the shortest path of the generated mazes returned by the best program and $D_{avg}$ is the average distance of all executable programs. Each program generates five mazes for evaluating $D_{max}$ and $D_{avg}$. S.Iter. is the percentage of iterations that resulted in working code.
Model $D_{max}$ $D_{avg}$ S.Iter.
Llama 3.1 8B 11.40 5.25 90%
Llama 3.1 70B 9.80 2.50 100%
Llama 3.1 405B 14.60 2.63 90%
Claude 3.5 Haiku 29.40 5.58 100%
Claude 3.5 Sonnet 16.40 4.90 100%
GPT 4o mini 29.40 11.82 100%
GPT 4o 29.40 11.07 100%
o1-mini 17.40 3.11 100%
Mistral Small 8.40 1.84 80%
Mistral Large 7.40 0.83 100%
Gemini Flash 28.40 8.51 70%
Gemini Pro 4.00 -0.08 100%
Table VI shows the results for the maze generation experiment. In contrast to the previous experiments, all small models beat their larger counterparts. The only large model that is on par with the smaller ones is GPT 4o, which loses slightly to GPT 4o mini regarding the average distance $D_{avg}$. o1-mini also falls behind both GPT 4o models. The worst model is Gemini Pro, which fails to even connect the start and end points of the maze most of the time, resulting in a negative average reward, while always producing executable programs. Even the smallest LLM, Llama 3.1 8B, which was one of the worst models in the previous experiments, is better than the larger Llama models in terms of $D_{avg}$ and only loses by a small margin to the largest Llama model in terms of $D_{max}$. It should be further investigated why larger LLMs have difficulties in generating good mazes, but this is beyond the scope of this paper.
V-E Python Code Evaluation
Table VII shows the summary statistics of the synthesized Python code for the previous experiments. The two smaller Llama models are provided for free by Google Cloud as they are currently in public preview. It is visible that the more expensive models lead to more executable programs. For the successful iterations, the difference between smaller and larger models is not that significant. Claude Haiku is only 3.85% below the best model, Claude Sonnet, but costs less than a third of the overall cost of the experiments. Being a small, cost-efficient model, it is impressive that Claude Haiku beats almost all of the larger models, falling only behind Claude Sonnet and OpenAI’s o1-mini. This was also evident in the previous experiments where Claude Haiku produced the best programs for Asterix, Space Invader, for $\omega=30$ in the vehicle driving experiment, for level 8 of Baba is you and was also on par with two other LLMs for the maze generation experiment. This indicates that large, expensive models are not always necessary and that, depending on the task, small models are just as good as expensive ones, especially when an evolutionary search strategy is used.
@@ V-E Python Code Evaluation
TABLE VII: The overall evaluation of the synthesized Python code. Cost is the total cost of all Python experiments. S.Iter. is the percentage of iterations that resulted in working code and returned a positive reward. Exec. Programs is the percentage of all generated programs that the Python interpreter could run.
Model Cost (\$) S.Iter Exec. Programs
Llama 3.1 8B Free 11.92% 57.95%
Llama 3.1 70B Free 20.00% 85.64%
Llama 3.1 405B 16.11 20.00% 82.31%
Claude 3.5 Haiku 5.19 26.92% 89.49%
Claude 3.5 Sonnet 17.58 30.77% 91.54%
GPT 4o mini 0.63 18.08% 86.28%
GPT 4o 9.25 26.15% 94.74%
o1-mini 25.73 30.38% 95.26%
Mistral Small 0.73 10.00% 63.21%
Mistral Large 7.36 20.00% 83.97%
Gemini Flash 0.28 15.00% 80.64%
Gemini Pro 7.80 17.31% 88.46%
V-F Tabletop Games Framework (TAG)

### v2

@@ V-B Programmatic Policies: Minatar
Claude Sonnet 133.36 97.97 87.09 69.18 70.19 72.03 75.75 66.71 69.61 65.67 80.76
Claude Haiku 101.0 119.55 92.98 71.7 79.92 67.01 62.01 74.83 69.19 75.37 81.36
Gemini Flash 171.79 109.88 81.12 85.61 69.4 74.91 76.53 74.07 67.49 74.02 88.48
Gemini Lite 169.27 106.47 91.18 72.45 62.3 73.63 57.69 101.62 100.37 68.42 90.34
Mistral Large 113.69 76.54 82.5 68.81 68.42 83.33 55.62 63.58 61.58 57.04 73.11
Mistral Small 144.33 90.16 71.24 100.13 57.9 60.84 68.51 58.6 56.92 56.53 76.52
o1 mini 123.64 101.49 86.57 77.88 83.49 64.92 79.07 71.62 57.42 71.16 81.73
GPT 4o 127.58 99.21 78.93 111.06 64.91 99.81 95.28 83.24 74.2 61.57 89.58
GPT 4o mini 144.86 134.72 104.54 111.67 89.83 94.26 84.0 96.41 95.25 96.99 105.25
Llama 3.3 70B 152.17 106.87 111.51 119.46 105.09 106.36 77.21 97.09 75.86 101.72 105.33
Llama 3.1 70B 135.31 96.67 92.75 72.29 74.85 62.51 62.15 64.76 58.73 62.49 78.25
For Breakout, the best programs are all able to locate the ball and also determine the direction in which the ball is moving. Claude Sonnet and o1 mini even check whether the ball bounces off walls. The best performing program of o1 mini simulates the movement of the ball until it is in the same row as the paddle.
In Asterix, the best performing programs with a reward over 10 prioritise reaching the gold while checking whether it is moving towards an enemy and, if so, avoiding the enemy. The mediocre performing models often prioritize the gold without checking if there are enemies nearby. Claude Sonnet even takes into account the trail of the enemy to check if the chosen action is safe.
In Space Invader, the good programs with a reward of over 20 correctly locate the enemy bullets, the aliens and the cannon and then use threat detection to check whether the enemy bullets need to be dodged before the enemies themselves are shot. The programs with a score below 20 do not anticipate the movement of aliens or prioritise shooting enemies over avoiding enemy bullets.
Overall, it can be said that in the Minatar games larger models on average show more sophisticated behaviour in the programs, but as can be seen with Claude Haiku or Gemini Flash Lite, this is not always the case, since for some games they achieve a better reward. Currently, only a very simple prompting strategy is used, which already gives comparable results to some of the baselines reported in [17] or even outperforms all baselines in the case of Breakout. Using more complicated prompting strategies, such as Chain of Thought [50] or adding a crossover operator could lead to improvements in the programs found.
V-C Vehicle Driving
@@
Table IV shows the results of the driving experiments for different rotation speeds $\omega$ to adjust the difficulty level of steering the asteroid ship. The numbers are the minimum distance achieved by the best program for five evaluation episodes. $D_{avg}$ is the average of all distances for each LLM. We omit the number of successful iterations because no LLM managed to stop the asteroid ship at the target position in all five evaluation episodes. A program was considered successful if it could consistently stop the vehicle within a specified tolerance $t$ across all evaluation episodes. In our experiments, the synthesized programs succeeded in stopping the vehicle in only one or two episodes within a tolerance of $t=10$, and thus no program qualified as successful. The best models overall are the two Mistral LLMs, which together achieve the best programs for 9 out of 10 rotation speeds. Only for $\omega=10$ did another LLM (Claude Haiku) find the best program. In terms of $D_{avg}$, larger models generally outperformed their smaller counterparts but usually only by a small margin. As no LLM successfully solved the problem with a simple prompting strategy in this experiment, we consider it a compelling challenge for future research.
@@
TABLE V: Highest reward per language model and level (with number of successful trials per level). The color of the heatmap is based on the successful trials, as it is more important to know how many times the LLMs were able to solve a level out of all 10 trials or if the trials were solved randomly.
@@
Claude Sonnet 95 (10) 89 (5) 0.0 (0) 95 (10) 0.0 (0) 91 (1) 94 (10) 95 (10) 90 (1) 92 (3) 8
Claude Haiku 95 (10) 0.0 (0) 0.0 (0) 95 (10) 0.0 (0) 0.0 (0) 94 (10) 95 (10) 0.0 (0) 0.0 (0) 4
Gemini Flash 95 (10) 0.0 (0) 0.0 (0) 95 (2) 0.0 (0) 0.0 (0) 94 (7) 95 (9) 0.0 (0) 0.0 (0) 4
Gemini Lite 95 (10) 0.0 (0) 0.0 (0) 95 (2) 0.0 (0) 0.0 (0) 94 (7) 95 (9) 0.0 (0) 0.0 (0) 4
Mistral Large 95 (6) 89 (1) 0.0 (0) 72 (1) 0.0 (0) 0.0 (0) 94 (9) 95 (7) 0.0 (0) 0.0 (0) 5
Mistral Small 95 (6) 0.0 (0) 0.0 (0) 63 (1) 0.0 (0) 0.0 (0) 94 (2) 95 (8) 64 (2) 0.0 (0) 5
o1 mini 95 (10) 89 (2) 0.0 (0) 95 (10) 0.0 (0) 0.0 (0) 94 (10) 95 (10) 82 (1) 92 (5) 7
GPT 4o 95 (10) 89 (1) 0.0 (0) 95 (9) 0.0 (0) 0.0 (0) 94 (8) 95 (9) 0.0 (0) 92 (1) 5
@@
Llama 3.3 70B 95 (10) 0.0 (0) 0.0 (0) 95 (3) 0.0 (0) 0.0 (0) 94 (6) 95 (6) 0.0 (0) 0.0 (0) 4
Llama 3.1 70B 95 (10) 0.0 (0) 0.0 (0) 95 (2) 0.0 (0) 51 (1) 94 (4) 95 (6) 0.0 (0) 0.0 (0) 5
V-D Baba is You
@@
For our experiments, we used a Python version [^11] https://github.com/ADockhorn/Keke-AI-PY of the Keke is You AI framework [18]. For this domain, we prompted the LLMs to provide a policy, giving a short description of the game and the initial state of the level (the complete prompt is in the repository). Similar to the Minatar experiments, the state is converted into a text description. The function to be written should use the current state as input and return a movement direction or the command for waiting a turn. Each episode ends after 100 actions or once the win condition is fulfilled. A reward is awarded based on the maximum number of actions (100) minus the number of steps taken. Thus, the return can be maximized by finishing the level as fast as possible. Each level can be solved in less than 20 actions.
In our tests, we queried the agent to solve 10 simple demo levels (see Figure 4). Each of the levels focuses on one or more key mechanics of the framework such as rule interpretation (levels 1-10), rule creation (levels 2, 3, 5) or destruction (levels 6, 8, 9, 10), and object manipulation (level 7). Table V shows the results of our comparison of the LLM models’ capabilities. The number of successful iterations is shown in brackets.
All agents were able to solve at least 3 out of 10 levels, with Claude 3.5 Sonnet being the only model able to solve 8. For Claude and GPT, models of the same vendor with a higher number of parameters were able to solve more levels. For the Llama models, the 3.1 version solved one more level than the 3.3 model. In case of Gemini and Mistral, both the small and large models performed about the same. Tested models were mostly successful in interpreting existing rules. As can be seen, some levels are rarely solved by any model. Creating or destroying rules and thus modifying the logic of our game world has proven difficult for all models. Many models failed in solving levels 2 and 3, which require rule creation, and levels 9 and 10, which require a rule’s destruction to finish the puzzle. Slight differences in the observed success rate could be due to the low number of repetitions per level, resulting in sampling errors: levels 2, 6, 9, and 10, which are rarely solved at all, could be affected by this. Chain of thought prompting [50] may help in overcoming these more complex planning tasks.
V-E Procedural Content Generation
Figure 5: Three generated mazes aiming to optimise the longest shortest path objective. Left: (score 18) example from a simple LLM generated algorithm setting wall cells with a fixed probability. Middle: (score 38) example a more sophisticated LLM algorithm involving recursion and a shortest path algorithm. Right: (score 54) example from an evolutionary algorithm directly optimising for the objective function.
@@ V-E Procedural Content Generation
The prompt advised the LLMs to use the longest shortest path objective to guide the maze generation process. This objective encourages intricate and interesting mazes. Most of the generated code ignored the hint and instead coded overly simple algorithms, placing corridors and walls in each cell with a given probability while usually ensuring that the start and end points were not on wall cells. An example generated maze is shown in the left of figure 5. Occasionally, a better algorithm was produced that mixed randomness, recursion and graph search in ways we have not fully analysed. These algorithms sometimes produced mazes with no path between start and end, resulting in a score of -1. When they worked, they often produced reasonable mazes such as the one shown in the middle of Figure 5. The LLMs failed to find an algorithm as effective as an evolutionary algorithm applied to directly solve the objective. A sample maze from such an algorithm is shown on the right of the figure.
@@
Note that here we are evaluating the effectiveness of the algorithms in meeting the specified objective, which is to produce mazes with the longest shortest path between start and end. Depending on the application, this could be a poor objective to maximise, with the best mazes having a mid-ranking score, such as the central maze in Figure 5.
@@
TABLE VI: Maze generation LLM results. $D_{max}$ is the maximum distance of the shortest path of the generated mazes returned by the best program of all trials and $D_{avg}\pm\sigma$ is the average distance with the standard deviation of the 10 programs found after all trials. Each program generates five mazes for evaluating $D_{max}$ and $D_{avg}\pm\sigma$.
Model $D_{max}$ $D_{avg}\pm\sigma$
Claude Sonnet 18.0 8.98$\pm$3.79
Claude Haiku 29.4 15.4$\pm$8.73
Gemini Flash 17.0 15.7$\pm$1.28
Gemini Lite 23.2 14.36$\pm$5.46
Mistral Large 10.8 4.2$\pm$3.3
Mistral Small 10.8 4.6$\pm$3.65
o1 mini 44.4 18.56$\pm$15.92
GPT 4o 33.4 18.48$\pm$11.66
GPT 4o mini 33.4 22.02$\pm$11.76
Llama 3.3 70B 14.8 8.3$\pm$3.75
Llama 3.1 70B 11.4 6.76$\pm$2.67
Table VI presents the results of the maze generation experiment. Unlike the previous tasks, smaller models perform in line with, or even outperform, their larger counterparts. Among the larger models, only GPT-4o achieves performance on par with the smaller models, although it performs slightly worse in terms of $D_{avg}$. Mistral Large also performs similarly to Mistral Small, but both models cannot keep up with models from other LLM providers. As with Table III, the high standard deviation of $D_{avg}$ in this experiment supports the additional long-running experiments in Section VI.
The reasoning model o1-mini outperforms both variants of GPT-4o with respect to $D_{max}$, but falls short compared to GPT-4o mini in terms of average distance. Overall, most models struggle with this task, with the exception of the OpenAI models and Claude Haiku. Understanding why larger language models underperform in this domain remains an open question, but is beyond the scope of this paper.
V-F Python Code Evaluation
Table VII shows the summary statistics of the synthesized Python code for the Minatar, Baba is You, maze generation and vehicle driving experiments, including the costs incurred. On average, the larger models do better than their small counterpart. In the case of Llama the newer 3.3 model beats the 3.1 model in terms of the successful iterations and trials. The successful iterations are, however, quite similar for all LLMs. The Mistral models have by far the worst percentage of executable programs but are still on the 4th and 6th rank in terms of overall performance, which indicates that the percentage of executable programs are not a good proxy for selecting an LLM to generate game-related Python code. o1 mini is on average the best model on all Python domains, and is also the only reasoning model in this benchmark. This indicates that reasoning models have an advantage in generating game-specific code. A more fine-grained ranking of LLMs for each domain is presented in Table X in Section VII, where we discuss the experiments in general and also give some practical recommendations.
@@ V-F Python Code Evaluation
TABLE VII: The overall evaluation of the synthesized Python code. Cost is the total cost of all 2600 iterations for the LLM (100 per game) (the Llama 3.1 70B model is provided for free by Google Cloud as it is currently in public preview). S.Iter. is the percentage of iterations that resulted in working code and S.Trl the percentage of trials (each consisting of 10 iterations). Only programs that returned a positive reward were considered for S.Iter and S.Trl. Exec. Programs is the percentage of all generated programs that the Python interpreter could run. For each of the four domains each LLM is ranked from 1 to 11 based on the performance of the best result for the domain. This rank is averaged and used to assign an overall Rank.
Model Cost (\$) S.Iter S.Trl Exec. Programs Rank
Claude Sonnet 80.75 63.27 80.77 83.23 2nd
Claude Haiku 17.34 61.35 76.92 86.35 5
Gemini Flash 2.56 62.62 72.31 95.85 7
Gemini Lite 1.30 61.15 71.54 94.28 10
Mistral Large 43.57 60.08 70.77 75.37 4
Mistral Small 2.06 59.65 68.85 74.7 6
o1 mini 35.87 62.88 80.0 90.93 1st
GPT 4o 36.11 62.19 76.15 92.09 3rd
GPT 4o mini 2.18 62.0 67.31 91.09 9
Llama 3.3 70B 6.86 62.08 71.15 92.34 11
Llama 3.1 70B 0.00 61.08 70.38 82.59 7
V-G Tabletop Games Framework (TAG)